OpenStudy (anonymous):

how to find x ,for log(x/2)=2 ?!

OpenStudy (anonymous):

10^log(x/2)=10^2 x/2=100 x=200 is it true ?

ganeshie8 (ganeshie8):

\( \log(x/2)=2 \) \( x/2=10^2 \) \(x = 2 \times 10^2\)

ganeshie8 (ganeshie8):

\(\large \color{red}{\checkmark}\)

OpenStudy (anonymous):

thanks

ganeshie8 (ganeshie8):

np :)